The cost of a rental truck to move furniture includes a fixed daily cost and a variable cost based on miles driven. For a customer driving 180 miles, the cost for 1 day will be $75. For a customer driving 150 miles the cost for 1 day will be $67.50. determine the fixed cost and the cost per mile for this truck.
How do you set this up.

Answer by richwmiller(17219)

the basic equation is a linear equation
y=mx+b
or
t=mr+db
75=180*r+db,
67.50=150*r+db
where d=1
75=180*r+b,
67.50=150*r+b
b = 30., r = 0.25
.25 cents a mile
and base rate of 30 dollars a day